OpenStudy (anonymous):

The weekly sales in thousands of items of a product has a seasonal sales record approximated by n = 84.59 + 24.8 sin πt/24 (t is time in weeks with t = 1 referring to the first week in the year). During which week(s) will the sales equal 96,990 items? A. week 4, week 20, and week 52 B. week 21 and week 30 C. week 4 and week 47 D. week 30 and week 47

OpenStudy (anonymous):

(96.990 - 84.59)/24.8 = sin (t)(pi)/24 = 0.5 This happens at t = 4, 20, and 52 because those values will correspond to: sin pi/6, sin 5(pi)/6, sin pi/6

OpenStudy (anonymous):

All good now, @Girl_Ninja ?

OpenStudy (anonymous):

oh okay yes @tcarroll010 thank you ;-)

OpenStudy (anonymous):

You're welcome! Good luck in all of your studies and thx for the recognition! @Girl_Ninja